OpenStudy (anonymous):

What is the DOMAIN & RANGE of : g(x)=(1/√ 4-(x^2))+3

OpenStudy (anonymous):

\[{1\div \sqrt{4-x^{2}}} +3\]

OpenStudy (anonymous):

is this your question?

OpenStudy (anonymous):

1/\[\frac{ 1 }{ \sqrt{4-x^2}} +3\]

OpenStudy (anonymous):

well domain is (2,-2)

OpenStudy (anonymous):

oops,it should be (-2,2)

OpenStudy (anonymous):

How do you know that?

OpenStudy (anonymous):

anything which is inside a square root must always be positive or equal to zero,so that it is real.do you know that?

OpenStudy (anonymous):

Oh okay so do I just set it to equal zero and solve...? But I thought that was finding the asymptotes

OpenStudy (anonymous):

i mean